OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1)with a slope of 4. Answer y - 3 = 4(x + 1) y - 1 = 4(x - 3) y + 1 = 4(x + 3) y + 3 = 4(x - 1)

OpenStudy (nenadmatematika):

y+1=4(x+3)

OpenStudy (amistre64):

given a point (Px,Py), and a slope "m" the equation of the line in point slope format would then be: y - Py = m(x - Px)